# An LLM-based Recommender System Environment

### v1

@@ 4.1 Setup
In this section we describe the standard configuration of our framework, we also conduce ablations experiments that justify these choices, for more details see Section 5.
We retrieves items from the history based on their similarity to the query item. We generate item-description embeddings using a Sentence-T5 model (Ni et al., 2022) and compute their similarities based on the cosine distance. To select the item-rating pair to retrieve from memory, we compute the similarity between the query item and all items previously viewed by the current user, selecting the items with the highest similarity. We use the items returned from the retrieval component to construct a prompt to query the LLM. The LLM is tasked with generating a rating of the queried item by the current user, where the queried item corresponds to the item suggested by the recommender system. We construct the prompt such that the user description comes first, allowing us to leverage the key-value cache (Pope et al., 2023), eliminating the need to recalculate all intermediate embeddings within the layers of the LLM for already encountered prefixes, therefore, increasing execution speed. We use few-shot prompting to improve model performance as shown by Brown et al. (2020). In addition we use a custom system prompt see Appendix A for movie and Appendix B for books. We experimented with various GPTQ (Frantar et al., 2022) versions of Llama, Vicuna, Mistral using LLMs that could run within a 24GB memory limit, for more details see Appendix C and we choose Mistral-7B as default.
Our implementation of reward shaping operates on the following premise: as a user engages with an item more frequently, their interest in revisiting it diminishes. Conversely, as time passes, the likelihood of the user interacting with the item increases again (Russell and Levy, 2011). Given this insight, let us consider a user $u$ from the set $U$ and an item $i$ with which the user has interacted $n_{ui}$ times. When a time span of $\Delta t$ has passed since the last interaction with the item, the reward $r$ undergoes a reshaping process, characterized by the following equation:
@@
Table 2: Results from RL methods trained on SUBER: “Liked genres” represents the percentage of user-preferred genres among the top ten recommendations (cf. Appendix E). “Pers.@10” denotes personalization within the initial ten recommendations.
Algorithm Average
reward
MAP@10 $\uparrow$ MMR@10 $\uparrow$ Pers.@10 $\uparrow$ % Liked
genres $\uparrow$
% Disliked
genres $\downarrow$
Table 3: Random examples of movie recommendation from trained A2C model on SUBER.
Watched movies by user Movies recommended to user
The Walk, The Testament of Dr. Mabuse The Island, Captain America, The Count of Monte Cristo
Inception, The Walk Spartacus, Exodus: Gods and Kings, Cliffhanger
4.2 Human Evaluation
We conducted a case study to better evaluate the quality of different LLMs in the rating simulation task. For the study, we sampled ten user-movie interactions, for each interaction we queried four different LLMs. The random rating in Table 4 serves as a baseline comparison, allowing us to compare the quality of our proposed approach against a random signal. The answer is constructed by sampling a rating uniformly at random between 1 and 10, and having the LLM (Vicuna 13B) generate the explanation for the rating. We then asked participants to rate the quality of the LLM’s response on a scale of 1 to 5. More information on the study setting is provided in Appendix D.
Table 4: Human evaluation scores for various LLMs. 14 participants were asked to rate the response of the LLM between 1 and 5. We observe that humans strongly prefer GPT-4’s answers compared to other LLMs.
LLM Score $\uparrow$
Random rating 2.87$\pm$1.51
Vicuna 13B 3.22$\pm$1.32
Llama-2-Chat 13B 3.42$\pm$1.22
Mistral 7B 3.80$\pm$1.27
GPT-4 4.47$\pm$0.77
4.3 Benchmarks
We demonstrate the viability of our environment to train an RL recommender system. The RL model architecture is inspired by the principles of Low-Rank Approximations in collaborative filtering (Aggarwal et al., 2016). We implemented four different agents based on A2C (Mnih et al., 2016), PPO (Schulman et al., 2017), TRPO (Schulman et al., 2015), and DQN (Mnih et al., 2013). We train all models for 1.6M steps on SUBER. Due to space constraints, a more detailed discussion on the training of the reinforcement learning models is deferred to Appendix E. Our evaluation indicates that the A2C algorithm demonstrates the best overall performance in our case study (cf. Table 2). While the PPO algorithm registers a higher personalization score, indicative of its ability to tailor recommendations, it is less effective than A2C in aligning recommendations with user interests, as reflected in the percentage of liked genres metric. This suggests that A2C is more adept at discerning and catering to user preferences. To provide a qualitative insight into the performance of our models, we present two examples of movie recommendations generated by the A2C algorithm (cf. Table 3).
5 Ablations
5.1 Method
@@
To determine the effect of each component in our environment, we performed ablations across four different test cases.
Genres/Categories
@@
We assess the environment’s ability to recognize movie and book genres, and its ability to correlate those genres with user preferences to accurately predict ratings. User profiles were created manually for each movie genre, ensuring that they express a preference for the selected genre while disliking all others. Afterwards, we queried the environment with users and movies from both their favored and disliked genres. The accuracy of rating predictions is used to measure performance. A similar process is used for the book environment, replacing movie genres with book categories.
High/Low Rating
Collection of Items.
Similarity to Real Rating Distribution
We evaluate whether the rating distribution obtained from our movie environment accurately reflects human behavior by comparing it to the rating distribution from MovieLens, which are representative samples of human ratings. We sample with replacement from both our environment and the MovieLens dataset. We then compute the empirical distribution across the dataset and use the total variation distance as a metric to measure similarity.
Aggregated Score
The aggregated score is the mean of all test cases. All ablations, except where defined otherwise, were performed using the following configurations. We used the 2-shot prompting, a custom system prompt, three item retrievial via T5-similarity, and no reward perturbation. For movies, we used Mistral 7B with rating scale 0-9, and for books we use Mistral 7B with scale 1-5.
5.2 Retrieval Component
We implemented three different approaches for the retrieval component: feature retrieval, recency retrieval, and similarity retrieval. The feature-based approach retrieves items based on the Sorensen Coefficient of movie genres, actors, director, and average rating, while for books, we use book category, author, and average rating. The recency algorithm retrieves the most recent interactions. The similarity approach retrieves items from the history based on their similarity to the query item. We generate item-description embeddings using a Sentence-T5 model (Ni et al., 2022) and compute their similarities based on the cosine distance. To select the item-rating pair to retrieve from memory, we compute the similarity between the query item and all items previously viewed by the current user, selecting the items with the highest similarity.
Our ablation of the retrieval component demonstrates that this component plays a crucial role in understanding user interests (cf. Table 5). Furthermore, the recency approach proves inadequate, while the best-performing retrieval approach is predicated on the similarity of item features.
Table 5: Comparison of different retrieval strategies on the collection of items test set for movies and books. We only show the entries of the test case where the result changed significantly, for extended results see Appendix C.
Retrieval
component
Collection of movies $\uparrow$ Collection of books $\uparrow$
None 0.49$\pm$0.01 0.50$\pm$0.00
Most recent 0.62$\pm$0.01 0.64$\pm$0.01
T5 similarity 0.67$\pm$0.03 0.76$\pm$0.02
Feature similarity 0.68$\pm$0.01 0.79$\pm$0.01
5.3 Prompting
Furthermore, we experimented with one-shot and two-shot prompting to improve model performance, which has been shown to increase generation quality (Brown et al., 2020). In addition to the default system prompt , we created a custom system prompt see Appendix A for movie and Appendix B for books. Tokenization ambiguity can become an issue when generating numbers with LLMs. Since we are dealing with ratings on a scale from one to ten, and because the number “10" can be tokenized in two different ways, this can cause unwanted side effects. To tackle this challenge, we tested two additional strategies for the movie setting: shifting all rewards to the scale of 0-9, and using words for numbers from “one" to “ten."
In the movie environment, we observe that different prompt strategies generally do not differ significantly from each other in the case of Mistral, with the only two exceptions being the 0-shot prompt with the default system prompt, which performs slightly worse, and the weak performance of the 1-10 rating scale due to tokenization ambiguity. Vicuna, on the other hand, is more affected by different prompt strategies, as shown in Eq. 4. Table 6 shows the general trend on how the environment is able to capture human concepts such as genres and movie franchises. For the book environment see Table 12.
@@ 5.3 Prompting
5.4 LLMs
Since different LLMs influence the simulation of human behavior in different ways, it is important to highlight the inherent trade-off between model size and processing speed. In particular, during training of an RL model, a fast environment is desirable to acquire more samples in a shorter time span. However, smaller LLMs may not adequately emulate the desired human behavior of our synthetic users.
In general, we observe that larger models perform better across model families (cf. Figure 3(a)). In addition, we can see how Mistral performs best among open-source models. In addition, it can be seen that the ranking of the aggregated scores of the different models and the ranking obtained with the human evaluation are largely consistent. In fact, GPT4 gets the best score, followed by Mistral 7B, while Llama-2-Chat 13B and Vicuna 13B get similar but slightly lower scores.
(a) Movies
(b) Books
@@ 5.4 LLMs
Figure 3: Aggregated score across LLM families for the movie environment (a), and for the book environment (b) by varying only the LLM component. For numerical values of the scores of each test case, see Appendix C.
@@ 6 Conclusion
Our research offers a possible avenue to address the persistent challenge of training recommender systems in the absence of real user interactions. Conventional approaches that depend on user-item interaction histories or synthetic data have often failed to replicate real-world usage scenarios accurately. By introducing SUBER, a novel RL environment designed specifically for recommender system training, and incorporating recent advances in LLMs to emulate human behavior in the training environment, we have proposed a potential solution to this long-standing issue. Our results, as demonstrated through a series of ablation studies, experiments and human evaluation, underscore the efficacy of our approach. We believe that this work marks a step toward achieving more realistic and practical training environments for recommender systems, even when direct user interactions are unavailable.
@@ A.1 Users Generation
When generating synthetic users, our process begins by randomly sampling an age from a distribution reflecting age demographics in the United States (Bureau, 2022). In addition, we randomly select a hobby and a profession from predefined lists. These hobby lists are divided into two categories: one tailored to children (aged 4-17) and another for adults (aged 18-75). Users not of working age are assigned the profession “student," while those of retirement age are categorized as “retired."
In total, children users can possess one of 33 hobbies, while adult users have a choice of 422 hobbies. Regarding professions, there are 200 different options available. Once all user attributes are determined, they are incorporated into a prompt that generates a user description (see LABEL:listing:prompt for an example). All lists are generated using the GPT-3.5 model, with the exception of the adult hobbies list, for which we utilized data from (Raj, 2022).
For a comprehensive list of generated hobbies and professions, please refer to Appendix F. An illustrative example of a synthetic user can be found in LABEL:listing:prompt.
To train the RL model, we created an additional dataset using a similar approach. The primary distinction lies in how we sampled the user’s preferred and disliked genres, which were not generated using the LLM. This modification was made to ensure that the dataset includes users with a more diverse range of preferences. We sampled the movie genres preference according to the distribution of preferred genres in the US in 2018 (Consult, 2018). We show in Figures 4(a) and 4(b) how this strategy affects the genre preference of users.
(a) Genre preferences of user generated via LLM.
(b) Real distribution of genre preferences.
Figure 4: For each movie genre, we show in blue the percentage of generated users who like the genre. Similarly, we show in red the percentage of users who do not like the genre. (a) shows the distribution when the users are generated by the LLM. (b) shows the distribution of genres sampled according to genre popularity in the US.
@@ A.2 Items
Table 8: For each movie we retrieve the features showed in the table, in our implementation we only use a subset to describe the item.
@@
We use the same set of movies as contained in MovieLens (ml-latest-small) (Harper and Konstan, 2015) for our experiments and collect the respective movie features from TMDB, we show which feature we use in Table 8.
@@ A.3 Prompting
Both the approach of using digits from 0 to 9 and the word-based approach are designed to address tokenization ambiguity. This ambiguity arises because the number 10 can be tokenized in two different ways: as the token “10" directly or as separate tokens “1" and‘ ‘0." We also explored generating numbers directly without restricting them to a single token in the 1-10 approach. However, this approach exhibited poor performance, leading us to refrain from further experimentation.
@@
Q: is a
@@
Listing 2: Example query for rating to the LLM (Vicuna-v1.5-13B) using rating scale one-ten.
@@ A.3.1 Custom System Prompt
We also experimented with various system prompts, which are predefined text or instructions used to initiate a conversation or request from a user when interacting with a language model. The primary objective was to encourage the model to generate ratings that are less biased and more closely aligned with the information provided to the model. This includes factors such as the user description, the list of movies watched previously, and the overview of the queried movie, all of which play a role in shaping the predictions of a model. In Listing 3, we present our customized system prompt utilized for various analyses in Section 4.
@@
Listing 3: An advanced system prompt guiding the model to provide personalized and unbiased movie ratings based on detailed user and movie data.

### v2

@@ 4.1 Setup
We implemented three different approaches for the retrieval component: feature retrieval, recency retrieval, and similarity retrieval. The feature-based approach retrieves items based on the Sorensen Coefficient of movie genres, actors, director, and average rating, while for books, we use book category, author, and average rating. The recency algorithm retrieves the most recent interactions. The similarity approach retrieves items from the history based on their similarity to the query item. We generate item-description embeddings using a Sentence-T5 model [29] and compute their similarities based on the cosine distance. To select the item-rating pair to retrieve from memory, we compute the similarity between the query item and all items previously viewed by the current user, selecting the items with the highest similarity. We use the items returned from the retrieval component to construct a prompt to query the LLM. The LLM is tasked with generating a rating of the queried item by the current user, where the queried item corresponds to the item suggested by the recommender system. We construct the prompt such that the user description comes first, allowing us to leverage the key-value cache [32], eliminating the need to recalculate all intermediate embeddings within the layers of the LLM for already encountered prefixes, therefore, increasing execution speed. Furthermore, we experimented with one-shot and two-shot prompting to improve model performance, which has been shown to increase generation quality [7]. In addition to the default system prompt , we created a custom system prompt (see LABEL:listing:our_system_prompt for movie and Appendix B for books).
Listing 2: An advanced system prompt guiding the model to provide personalized and unbiased movie ratings.
Tokenization ambiguity can become an issue when generating numbers with LLMs. Since we are dealing with ratings on a scale from one to ten, and because the number “10" can be tokenized in two different ways, this can cause unwanted side effects. To tackle this challenge, we tested two additional strategies for the movie setting: shifting all rewards to the scale of 0-9, and using words for numbers from “one" to “ten."
We experimented with various quantized versions of Llama, Vicuna, Mistral, using LLMs that could run within a 24GB memory limit. A list of the models used in our experiments can be found in Appendix D. All models were quantized using GPTQ. Since different LLMs influence the simulation of human behavior in different ways, it is important to highlight the inherent trade-off between model size and processing speed. In particular, during training of an RL model, a fast environment is desirable to acquire more samples in a shorter time span. However, smaller LLMs may not adequately emulate the desired human behavior of our synthetic users.
For the reward perturbation experiment, we compared Gaussian noise and greedy noise. Greedy noise alters the LLM rating by 1 with a probability of $q$, while it remains unchanged with a probability of $1-q$.
Our implementation of reward shaping operates on the following premise: as a user engages with an item more frequently, their interest in revisiting it diminishes. In contrast, as time passes, the likelihood that the user interacts with the item increases again [38]. Given this insight, let us consider a user $u$ from the set $U$ and an item $i$ with which the user has interacted $n_{ui}$ times. When a time span of $\Delta t$ has passed since the last interaction with the item, the reward $r$ undergoes a reshaping process, characterized by the following equation:
@@
4.2 Ablations
@@
To determine the effect of each component in our environment, we performed ablations across four different test cases. In this section, we present the high-level idea; for more details, see Appendix C.
Genres/Categories.
@@
We assess the environment’s ability to recognize movie and book genres and its ability to correlate those genres with user preferences to accurately predict ratings. User profiles were manually created for each movie genre, ensuring that they expressed a preference for the selected genre while disliking all others. Afterwards, we queried the environment with users and movies from both their favored and disliked genres. The accuracy of rating predictions is used to measure performance. A similar process is used for the book environment, replacing movie genres with book categories.
@@
High/Low Rating.
Collection of Items.
Similarity to Real Rating Distribution.
We evaluate whether the rating distribution obtained from our movie environment accurately reflects human behavior by comparing it to the rating distribution from MovieLens, which are representative samples of human ratings. We sample with replacement from both our environment and the MovieLens dataset. We then compute the empirical distribution across the dataset and use the total variation distance as a metric to measure similarity. For the book environment, see Appendix C. The aggregated score is the mean of all test cases. All ablations, except where defined otherwise, were performed using the following configurations. We used the 2-shot prompting, a custom system prompt, three item retrievial via T5-similarity, and no reward perturbation. For movies, we used Mistral 7B with rating scale 0-9, and for books we use Mistral 7B with scale 1-5.
@@
Figure 3: Aggregated score across LLM families for the movie environment (top), and for the book environment (bottom) by varying only the LLM component. For details see Appendix D.
Results.
In the movie environment, we observe that different prompt strategies generally do not differ significantly from each other in the case of Mistral, with the only two exceptions being the 0-shot prompt with the default system prompt, which performs slightly worse, and the weak performance of the 1-10 rating scale due to tokenization ambiguity. Vicuna, on the other hand, is more affected by different prompt strategies, as shown in Appendix D . Table 2 shows the general trend on how the environment can capture human concepts such as genres and movie franchises. For the book environment (cf. Table 3) it can be observed that the use of few-shot prompting, as well as the custom system prompt, has a positive impact on the different test cases. Additionally, similar to the movie environment, the model is also able to understand human concepts in the book domain. In general, we observe that larger models perform better across model families (cf. Figure 3). In addition, we can see how Mistral performs best among open-source models.
Our ablation of the retrieval component demonstrates that this component plays a crucial role in understanding user interests (cf. Tables 8 and 12 in Appendix D) . Furthermore, the recency approach proves inadequate, while the best-performing retrieval approach is predicated on the similarity of item features.
Table 4: Performance Metrics of RL Models Trained on SUBER: Mean Average Precision (MAP@10), Mean Reciprocal Rank (MRR@10), Personalization of the top ten recommendations (Pers.@10). “Liked genres” indicates the proportion of movies in the top ten recommendations aligned with user-preferred genres (see Appendix F for details).
Algorithm Average
reward
MAP@10 $\uparrow$ MRR@10 $\uparrow$ Pers.@10 $\uparrow$ % Liked
genres $\uparrow$
% Disliked
genres $\downarrow$
4.3 Human Evaluation
Table 5: Human evaluation scores for various LLMs.
LLM Score $\uparrow$
Random rating 2.87$\pm$1.51
Vicuna 13B 3.22$\pm$1.32
Llama-2-Chat 13B 3.42$\pm$1.22
Mistral 7B 3.80$\pm$1.27
GPT-4 4.47$\pm$0.77
We conducted a case study to better evaluate the quality of different LLMs in the rating simulation task. For the study, we sampled ten user-movie interactions, for each interaction we queried four different LLMs. The random rating in Table 5 serves as a baseline comparison, allowing us to compare the quality of our proposed approach with a random signal. The answer is constructed by sampling a rating uniformly at random between 1 and 10, and having the LLM (Vicuna 13B) generate the explanation for the rating. We then asked participants to rate the quality of the LLM’s response on a scale of 1 to 5.
Participants in this study were recruited from among our colleagues and provided informed consent to participate. The study was designed with strict adherence to randomized double-blind procedures to ensure impartiality and reliability of the results. As this user study did not involve ongoing follow-up or monitoring of the participants, our institutional review board (IRB) determined that formal approval was not required. From the survey (cf. Table 5), we find that users agree more with GPT-4, outperforming all other models. Furthermore, we find that Mistral 7B is the best LLM among open-source models despite only having 7B parameters. More information on the study setting is provided in Appendix E.
4.4 Benchmarks
We demonstrate the viability of our environment to train an RL recommender system. The architecture of the RL model is inspired by the principles of Low-Rank Approximations in collaborative filtering [2]. We implemented four different agents based on A2C [27], PPO [40], TRPO [39], and DQN [26]. We train all models for 1.6M steps on SUBER. Due to space constraints, a more detailed discussion on the training of reinforcement learning models is deferred to Appendix F. In addition to using classical RecSys metrics, like $\text{MAP}@10$, $\text{MRR}@10$, and personalization, we introduce two additional metrics to assess the alignment of the agent’s recommendations with user preferences (See Appendix F for detail metric definitions). Each user in the training dataset has both preferred and disliked movie genres. Based on these data, the trained RL model generates a list of top-5 movie recommendations for each user: percentage liked genres and percentage disliked genres . The recommendations are classified into three categories: liked (movies matching preferred genres and excluding disliked ones), disliked (movies with disliked genres and without preferred ones), and neutral (remaining recommendations).
Our evaluation indicates that the A2C algorithm demonstrates the best overall performance in our case study (cf. Table 4). Although the PPO algorithm registers a higher personalization score, indicative of its ability to tailor recommendations, it is less effective than A2C in aligning recommendations with user interests, as reflected in the percentage of liked genres metric. This suggests that A2C is more adept at discerning and catering to user preferences.
5 Future Work
One promising direction is to fine-tune the LLM with human feedback to improve the simulated user behavior. This can be achieved using datasets like MovieLens, which provide a natural reward function for RL methods. For instance, the negative squared difference between the LLM rating and the actual rating can be used as a reward. Currently, the setup considers only static users. Future work could model user evolution over time to reflect changing interests, making synthetic users more realistic and dynamic. Additionally, exploring ways to enrich the feature space of the LLM could be valuable. By incorporating complex features such as item seasonality and user context, RL models could better capture user behavior, leading to more accurate simulations.
@@ 6 Conclusion
Our research offers a possible avenue to address the persistent challenge of training recommender systems in the absence of real user interactions. Conventional approaches that depend on user-item interaction histories or synthetic data have often failed to replicate real-world usage scenarios accurately. By introducing SUBER, a novel RL environment designed specifically for recommender system training, and incorporating recent advances in LLMs to emulate human behavior in the training environment, we have proposed a potential solution to this long-standing issue. Our results, as demonstrated through a series of ablation studies, experiments, and human evaluation, underscore the efficacy of our approach. We believe that this work marks a step toward achieving more realistic and practical training environments for recommender systems, even when direct user interactions are unavailable.
@@ A.1 Users Generation
When generating synthetic users, our process begins by randomly sampling an age from a distribution reflecting age demographics in the United States [8]. In addition, we randomly select a hobby and a profession from predefined lists. These hobby lists are divided into two categories: one tailored to children (aged 4-17) and another for adults (aged 18-75). Users not of working age are assigned the profession “student," while those of retirement age are categorized as “retired."
In total, children users can have one of 33 hobbies, while adult users have a choice of 422 hobbies. With regard to professions, there are 200 different options available. Once all user attributes are determined, they are incorporated into a prompt that generates a user description (see LABEL:listing:prompt for an example). All lists are generated using the GPT-3.5 model, with the exception of the list of adult hobbies, for which we used data from [35].
For a complete list of hobbies and professions generated, refer to Appendix G. An illustrative example of a synthetic user can be found in LABEL:listing:prompt.
To train the RL model, we created an additional dataset using a similar approach. The primary distinction lies in how we sampled the user’s preferred and disliked genres, which were not generated using the LLM. This modification was made to ensure that the dataset includes users with a more diverse range of preferences. We sampled the preference for movie genres according to the distribution of preferred genres in the US in 2018 [11]. We show in Figures 5 and 6 how this strategy affects the preference of users for the genre.
Figure 5: Genre preferences of user generated via LLM. For each movie genre, we show in blue the percentage of generated users who like the genre. Similarly, we show in red the percentage of users who do not like the genre.
Figure 6: Real distribution of genre preferences. For each movie genre, we show in blue the percentage of generated users who like the genre. Similarly, we show in red the percentage of users who do not like the genre.
@@ A.2 Items
We use the same set of movies as contained in MovieLens (ml-latest-small) [16] for our experiments and collect the respective movie features from TMDB, and show which features we use in Table 6.
Table 6: For each movie we retrieve the features shown in the table, in our implementation we use only a subset to describe an item.
@@ A.3 Prompting
Both the approach of using digits from 0 to 9 and the word-based approach are designed to address tokenization ambiguity. This ambiguity arises because the number 10 can be tokenized in two different ways: as the token “10" directly or as separate tokens “1" and‘ ‘0." We also explored generating numbers directly without restricting them to a single token in the 1-10 approach. However, this approach exhibited poor performance, which led us to refrain from further experimentation.
@@
Emily has previously watched the following movies (in parentheses are the ratings she gave on a scale of 0 to 9): "Twelve Monkeys" (7), "Star Wars" (6), "Top Gun" (5).
-romance
-drama
Here are the 2 main actors of the movie, in order of importance: Rade SerbedZija (M), Aleksandra Vujcic (F). On average, people rate the movie "Broken English" 5.6 on a scale of 0 to 9. Emily watches the movie "Broken English" for the 1st time.
What can you conclude about Emilys rating for the movie "Broken English" on a scale of 0 to 9, where 0 represents a low rating and 9 represents a high rating, based on available information and logical reasoning?
A: Based on Emilys preferences and tastes, I conclude that she will assign a rating of 7
Listing 3: Example query for rating to the LLM (Vicuna-v1.5-13B) using the 0-9 scale.
@@
Listing 4: Example query for rating to the LLM (Vicuna-v1.5-13B) using rating scale one-ten.
@@ A.3.1 Custom System Prompt
We also experimented with various system prompts, which are predefined text or instructions used to initiate a conversation or request from a user when interacting with a language model. The primary objective is to encourage the model to generate ratings that are less biased and more closely aligned with the information provided to the model. This includes factors such as the user description, the list of movies watched previously, and the overview of the queried movie, all of which play a role in shaping the predictions of a model. In LABEL:listing:our_system_prompt_app, we present our customized system prompt utilized for various analyses in Section 4.
@@
Listing 5: An advanced system prompt guiding the model to provide personalized and unbiased movie ratings based on detailed user and movie data.